OpenStudy (anonymous):

What are the x-intercepts of f(x)= 8/9(x^2+11x-80)

OpenStudy (anonymous):

for x-intercepts put f(x)=0 (i.e if y=f(x) we are using y=0) thus we have x^2+11x-80=0 and then solve using quadratic equations formula to get the required intercepts

OpenStudy (anonymous):

factorable.

OpenStudy (anonymous):

x^2+11x-80=0 x^2+16x-5x-80=0 or x(x+16)-5(x+16)=0 or (x+16)(x-5)=0 thus x=-16 or 5 thus the graph of f(x) crosses x-axis at -16,0 and 5,0